Caffeine (C 8 H 10 N 4 O 2 ), is a stimulant found in coffee, tea, colas and energy drinks. It is

used to restore mental alertness or wakefulness during fatigue or drowsiness. Caffeine is

also found in some headache and migraine medications. Solve for its percent


composition.


1
Expert's answer
2022-01-30T07:56:52-0500

Solution:

Atomic mass of elements:

Ar(C) = 12.0107 amu

Ar(H) = 1.00784 amu

Ar(N) = 14.0067 amu

Ar(O) = 15.999 amu


Calculate the molar mass of caffeine (C8H10N4O2):

M(C8H10N4O2) = 8×Ar(C) + 10×Ar(H) + 4×Ar(N) + 2×Ar(O)

M(C8H10N4O2) = 8×12.0107 + 10×1.00784 + 4×14.0067 + 2×15.999 = 194.1888 (g/mol)


Calculate the percent composition of caffeine (C8H10N4O2):

w(C) = [8×Ar(C) / M(C8H10N4O2)] × 100%

w(C) = [8×12.0107 / 194.1888] × 100% = 49.48%

w(C) = 49.48%


w(H) = [10×Ar(H) / M(C8H10N4O2)] × 100%

w(H) = [10×1.00784 / 194.1888] × 100% = 5.19%

w(H) = 5.19%


w(N) = [4×Ar(N) / M(C8H10N4O2)] × 100%

w(N) = [4×14.0067 / 194.1888] × 100% = 28.85%

w(N) = 28.85%


w(O) = [2×Ar(O) / M(C8H10N4O2)] × 100%

w(O) = [2×15.999 / 194.1888] × 100% = 16.48%

w(O) = 16.48%


Answer:

Caffeine contains 49.48% Carbon, 5.19% Hydrogen, 28.85% nitrogen, 16.48% oxygen.